### v2

# Local Structural Sisorder in Crystalline Materials

Abstract
ABSTRACT: Local positional disorder in soft, anharmonic materials has emerged as a central factor in shaping their electronic, vibrational, optical, and transport properties. Viewed mainly as a source of performance degradation, recent theoretical insights reveal that local disorder profoundly influences the electronic structure and phonon dynamics, without inducing deep electronic traps or non-radiative recombination pathways. In this Perspective, we highlight advances in modeling local disorder using polymorphous and anharmonic frameworks, showing how these methods explain experimental observations and predict new trends. We emphasize the role of disorder in the breakdown of the phonon quasiparticle picture and in modulating electron-phonon and phonon-phonon interactions, particularly in soft, anharmonic phases of matter, with significant effects on electrical and thermal transport. We outline opportunities for integrating these insights into predictive modeling for energy materials and propose combining advanced first-principles methods with machine learning.
Local disorder is an intrinsic feature of soft, anharmonic materials, first realized in the late 60’s by Comes et al. for oxide perovskites 1. While standard diffraction analyses of crystalline materials suggest high-symmetry monomorphous structures, local fluctuations in atomic positions give rise to a rich configurational landscape, as suggested by more advanced experimental approaches 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. These variations, linked to configurational entropy, are termed positional polymorphism 13 and impact key physical properties in ways that are not captured by conventional, monomorphous models. Some communities refer to this effect as static disorder or split-site occupancy, terms that are related but not strictly identical. Recent advances in first-principles theory have enabled the explicit treatment of such disorder 13, 14, offering a new lens to interpret and predict material behavior in energy applications.
Local disorder is a typical characteristic of room temperature phases of materials, and particular of anharmonic materials featuring a multiwell potential energy surface (PES). The high symmetry configuration corresponds to a local maximum on the PES, and locally disordered configurations correspond to energetically favorable minima, as shown in Figure 1a. At $T=0$ K, both cases correspond to static equilibrium states. At higher temperatures, ultraslow relaxation occurs between locally disordered configurations. The polymorphous framework does not explicitly model these dynamics, but instead represents structural disorder through optimized supercells or a limited sampling of them. This raises a fundamental question: which structural configurations should be considered as physically realistic for ab initio electronic structure calculations, particularly those based on density functional theory (DFT)? This choice also has profound implications for the coupling of carriers with lattice dynamics, i.e. electron-phonon coupling 15, as well as for phonon-phonon scattering which determine key optical and transport properties that govern the performance of energy devices.
In this Perspective, we present compelling evidence for the necessity of using locally disordered structures to achieve accurate predictions of electronic structure, vibrational dynamics, and electron-phonon interactions. Such an approach is crucial for capturing temperature-dependent properties, including vibrational spectra, effective masses, band gaps, and carrier mobilities, where theory results demonstrate excellent agreement with experimental data. We provide an analysis of ongoing theoretical and experimental efforts addressing the role of local disorder in materials, as well as outline limitations of the framework and emerging future directions. We discuss how local disorder can lead to improved predictions of key phenomena such carrier and thermal transport as well as polaron formation. We believe this Perspective will contribute to a deeper understanding of the intricate interplay between structural disorder, electronic structure, anharmonicity, and electron-phonon coupling, while also accelerating the development of more accurate theoretical frameworks for predicting energy-relevant properties of functional materials. Among the numerous materials known to feature local disorder 1, 16, 17, 18, 19, 20, 10, 21, emphasis is given to halide perovskites 2, 13, 22, a rapidly advancing area in energy research.
Figure 1: Impact of local disorder on the electronic structure. (a) Schematic representation of the PES illustrating the high symmetry and locally disordered structures, and the corresponding energy lowering $\Delta U$. (b) Locally disordered (polymorphous) structures of cubic CsPbI3 and CsSnI3 obtained by the ASDM using a $4\times 4\times 4$ supercell and DFT geometry optimization. (c) Average DFT charge densities of high symmetry and locally disordered cubic CsSnI3. (d) DFT electron spectral function of polymorphous CsSnI3 calculated by band structure unfolding 23. The band structure of the monomorphous model is shown as red. The band gap opening due to local disorder is $\Delta E_{\rm g}=0.56$ eV. Data are from Refs. 14, 24
1 Electronic structure
Local disorder in materials refers to spatially correlated deviations of atoms from their high symmetry positions, which on average preserve the crystallographic space group symmetry of the structure 13. To simulate positional polymorphism, a natural approach is to employ ab initio molecular dynamics (MD) with extended equilibration and production times to sample the relevant configurational space. Another approach is to start from a high symmetry supercell structure and introduce atomic displacements via random nudges 13 or special displacements along phonons 23, 14. The structure is then relaxed with the lattice constants fixed. This method enables the system to explore symmetry-breaking domains and identify energetically favorable minima in the PES, as illustrated in Figure 1a. The energy lowering with respect to the monomorphous structure is indicated as $\Delta U$.
[…]
Focusing on non-polar fluctuations, the degree of local disorder is quantified by calculating the deviation of the B-X-B angle ($\theta_{\rm B-X-B}$) from the ideal value of 180 in the cubic structure 24. In the examples shown in Figure 1b, local disorder is greater in CsPbI3 ($\Delta\theta_{\rm B-X-B}$ = 26.7) than in CsSnI3 ($\Delta\theta_{\rm B-X-B}$ = 23.2), which is also reflected in the corresponding $\Delta U$ values for each structure. Due to stronger orbital overlap and enhanced bond covalency, CsPbI3 has a larger lattice constant (6.25 Å vs. 6.14 Å in CsSnI3), which in turn provides greater structural flexibility and a higher tendency toward local symmetry breaking. As discussed in Ref. 22, another key parameter determining the degree of local disorder in cubic halide perovskites is the tolerance factor. This is strongly influenced by the size of the A-site cation, with higher values indicating less local distortions. For example, FA-based compounds have been shown to exhibit the lowest degree of positional polymorphism due to larger tolerance factors than MA-based and Cs-based compounds 24. Aziridinium (AZ)-based halide perovskites, a recent addition to the family of A-site cations 26, 27, are expected to exhibit a degree of local disorder in the range ($\Delta\theta_{\rm B-X-B}$ = 13–18) higher than FA-based compounds ($\Delta\theta_{\rm B-X-B}$ = 9–12), yet slightly lower than that typically observed in MA-based systems ($\Delta\theta_{\rm B-X-B}$ = 14–1922. This is because the estimated ionic radius of AZ is approximately 2.30 Å, which is intermediate between FA (2.53 Å) and MA (2.17 Å); the degree of local disorder in AZ-based compounds remains to be confirmed by first-principles calculations.
Figure 1c shows that using a monomorphous model in DFT calculations for cubic CsSnI3 and CsPbI3, including the effect of spin-orbit coupling (SOC), leads to an unphysical result. The charge density at the valence band maximum (VBM) and conduction band minimum (CBM) exhibits a spurious exchange in orbital character, with the VBM formed primarily by B-site metal p orbitals and the CBM showing contributions from X-site halogen p states and B-site s states. This results in a spurious negative gap of -0.27 eV and negative electron effective masses at the R-point for CsSnI3 as shown in Figure 1d. In contrast to common beliefs, this is not only a deficiency of DFT to handle electron correlations but a consequence of using the incorrect structural model without local disorder. As shown in Figure 1c for the average polymorphous structure, the expected orbital character in the charge density at the band extrema is recovered, leading to an overall band gap opening of 0.56 eV (Figure 1d). This value correlates with the B-X-B bond angle changes associated with local disorder and tolerance factor, consistent with trends calculated for other halide perovskites 22. Accounting for local symmetry breaking is essential to recover reliable band gap behavior, as also shown for oxide perovskites 17, low-dimensional quantum materials 20, transition metal monoxides 16, transition metal chalcogenides 28, and Cu-based superionic conductors 19. We stress that while DFT combined with locally disordered structures captures key qualitative features, accurate band gaps and effective masses require advanced methods like hybrid functionals or $GW$ to account more accurately for electron correlation 29, 24.
2 Structural and Vibrational properties
Figure 2: Impact of local disorder on structural and vibrational properties. (a) Calculated XRD patterns of monomorphous and polymorphous cubic SrTiO3 compared to experiment. Reproduced from Ref. 18, copyright (2021), with permission from Elsevier. (b) Calculated PDF of monomorphous and polymorphous cubic MAPbI3 compared with experiment. Reproduced with permission from Ref. 13, copyright (2020) by the American Physical Society. (c) Phonon spectral function of polymorphous CsPbI3 obtained by phonon unfolding. Data extracted from Ref. 14. (d) Phonon energies in reciprocal space measured for MAPbBr3 with inelastic neutron scattering. Figure adapted from Ref. 30 under $C$ BY 4.0. (e) Calculated Raman spectra of CsPbBr3 by MD compared to experimental measurements. Reproduced with permission from Ref. 31, copyright (2017) by the American Physical Society. (f) Calculated phonon DOS of monomorphous and polymorphous cubic Cu2Se compared with experiment. Data from Ref. 19.
[…]
Positional polymorphism also strongly affects the phonon spectra, leading to strongly overdamped vibrational dynamics, as shown for CsPbI3 by Zacharias et al. 14 (color map in Figure 2c). Employing a monomorphous structure together with the harmonic approximation to compute phonons results in pronounced dynamical instabilities in the phonon dispersion, as shown by the black lines in Figure 2c. This occurs because the monomorphous configuration corresponds to a local maximum on the PES, leading to negative curvatures along certain vibrational modes, manifested as imaginary phonon frequencies. Instead, using a polymorphous network yields dynamically stable phonons, with acoustic branches emerging near the $\Gamma$-point from a background of nearly dispersionless optical vibrations, consistent with inelastic neutron scattering measurements in halide perovskites 30, shown in Figure 2d for MAPbBr3. The strongly overdamped vibrations lead to the breakdown of the conventional phonon quasiparticle picture that is typically assumed in the monomorphous model. This breakdown becomes even more pronounced in hybrid halide perovskites due to enhanced mode mixing between polar optical phonons and the internal vibrations of the MA or FA molecules 24. We remark that in many systems, unstable modes become anharmonically stabilised at higher temperatures, still retaining a well-defined quasiparticle picture as shown by MD-based TDEP calculations 32.
Raman scattering measurements on halide perovskites reported by Yaffe et al. 31 reveal strongly correlated vibrational dynamics, as evidenced by a pronounced central peak spanning a broad frequency range, as shown in Figure 2e. This feature is absent in monomorphous harmonic systems, which typically exhibit sharp and well-defined Raman peaks. Interestingly, MD can reproduce the central peak of the high temperature phase of CsPbBr3, demonstrating that MD can capture both thermal vibrations and local disorder effects. Furthermore, MD has been proven powerful to capture overdamped phonon spectra in halide perovskites by analysis of velocity autocorrelation functions 33. Another observable that can be used to identify strongly overdamped, anharmonic vibrations is the phonon density of states (DOS). Figure 2e shows the phonon DOS of Cu2Se, a polymorphous material with high degree of local disorder, as calculated by Wang et al. 19 using the anharmonic special displacement method (ASDM) 23, 34. The presence of positional polymorphism leads to significant spectral broadening, in good agreement with experiment 35, and contrasts with the well-defined peaks calculated for the monomorphous structure. This behavior is consistent with prior observations of strongly overdamped vibrational spectra in superionic materials 36. Oxide perovskites represent another class of materials where overdamped phonon dynamics have been measured and supported through MD simulations 37.
The breakdown in the phonon quasiparticle picture, apart from resulting in broadened phonon spectra and reduced lifetimes, it also leads to enhanced phonon-phonon scattering, suppressing lattice thermal conductivity due to increased energy dissipation and reduced phonon mean free paths 38, 36, 19, 10. For thermoelectric materials, this suppression of thermal conductivity is beneficial, as it enhances the figure of merit, provided that the electronic transport remains relatively unaffected. Therefore, understanding and controlling disorder-induced phonon scattering is not only crucial for accurately modeling thermal transport but also offers a strategy to engineer high-performance thermoelectrics.
3 Thermal fluctuations in an anharmonic lattice
In Figures 3a and b, we illustrate the two sources of anharmonicity that influence the vibrational dynamics: one is positional polymorphism, as discussed for Figure 2, and the other arises from temperature-dependent phonon self-energy corrections. In the context of lattice dynamics, the loop diagram originating from quartic anharmonicity (Figure 3b) represents the first correction to phonon frequencies beyond the harmonic approximation 39. It corresponds to a process where a phonon interacts with the thermal fluctuations of the lattice, mediated by a virtual phonon through a quartic vertex. Although quartic terms are associated with the 4th-order potential, the loop diagram does not represent a real scattering process among four phonons. Instead, it captures a self-interaction mechanism, where thermal fluctuations in the lattice feed back into the mode’s own dynamics. This correction affects only the real part of the phonon self-energy and introduces no temperature-dependent broadening or lifetime effects, usually obtained by the so called bubble diagram 39. Within the self-consistent phonon (SCP) theory, temperature-dependent phonons are computed by iteratively solving an effective Hessian derived from interatomic forces. To obtain the final renormalized phonon frequencies and eigenvectors, SCP theory effectively resums anharmonic loop contributions to infinite order, offering a physically grounded and non-perturbative description of strong anharmonicity.
There are several state-of-the-art methods and codes that implement the SCP theory and extensions beyond it, including the Temperature Dependent Effective Potential (TDEP40, ALAMODE 39, stochastic self-consistent harmonic approximation (SSCHA41, and the ASDM 34. Within the same level of approximation, such as consistent calculation settings and order of anharmonicity, these methods are expected to yield identical results. We emphasize that these approaches to temperature-dependent anharmonicity are conceptually different than the polymorphous framework. They primarily incorporate phonon self-energy corrections either using thermally displaced configurations generated by MD, Monte Carlo, or special displacements sampling, yielding renormalized frequencies and linewidths, while preserving the average crystal symmetry and therefore do not explicitly explore local symmetry breaking. A key strength of these methods is the systematic use of crystal symmetries, which accelerates convergence and enables extensions beyond the loop diagram, for example, to capture corrections involving third- and higher-order anharmonicity 42, 32. In contrast, the polymorphous framework focuses on explicitly capturing local symmetry breaking and positional polymorphism. Phonon self energy corrections can be incorporated on top of the polymorphous framework starting from optimized supercells 14.
Figure 3: Impact of local disorder on electron-phonon coupling. (a,b,c) Locally disordered structure and diagrammatic representation of loop diagram contributing to temperature-dependent anharmonicity and the basic electron-phonon coupling process. Lines represent electrons with wavector $\bf k$} and $\bf k+q$} and the zigzag lines the anharmonic phonon. (d,e) Calculated band gap renormalization of monomorphous and polymorphous cubic CsPbI3 and tetragonal/cubic CsPbBr3 by ASDM compared with experiment. Data taken from Refs. 24, 14. (f) Calculated electron and hole effective mass of monomorphous and polymorphous cubic CsPbCl3. Data extracted from Ref. 24. Vertical dashed lines indicate temperatures where phase transitions take place.
4 Electron-phonon coupling
The anharmonic phonons obtained by accounting for local disorder, phonon self-energy corrections, or both can be coupled with the electronic structure of the locally disordered system, yielding the so called anharmonic electron-phonon coupling (Figure 3c). Electron-phonon coupling, in turn, affects both phonon and electron energies and lifetimes through self-energy corrections, as discussed extensively in Ref. 15. Figure 3d shows the effect of electron-phonon coupling on the band gap renormalization of cubic CsPbI3, as reported in Ref. 24. DFT calculations are performed within the ASDM using a $\times4$$\times$4 supercell. Results are averaged over 10 polymorphous configurations, weighted by their Boltzmann factors derived from free energies, thereby incorporating configurational entropy effects into the electron-phonon calculations. Employing polymorphous structures yields a positive band gap temperature coefficient, i.e., a band gap opening with increasing temperature. In contrast, calculations based on a monomorphous structure, not only underestimate the band gap, but also predict a band gap closing, due to an incomplete treatment of the Fan-Migdal self-energy contribution 22. In both cases, band gap renormalization is dominated by low energy optical modes (<<10 meV) with bending or rocking character, linked to B-X-B bond angle variations and octahedral tilting 22.
[…]
Figure 4: Impact of local disorder on transport and polaron formation. (a) Calculated electron mobility of monomorphous and polymorphous cubic SrTiO3 compared with experiment. Reproduced with permission from Ref. 43, copyright (2024) by the American Physical Society. (b) Calculated average charge-carrier mobility of orthorhombic MAPbI3 and cubic CsPbI3 compated with experiment. Reproduced with permission from Ref. 44, copyright (2019) by the American Chemical Society. (c) Polaron charge density in CsPbBr3 calculated by PBE0 in a supercell after adding a negative charge. Figure adapted from Ref. 45, copyright (2017) by the American Association for the Advancement of Science.
Figure 4a compares calculated electron mobilities, obtained using monomorphous and polymorphous structures of SrTiO3, with experimental data as reported in Ref. 43. The mobilities are computed using perturbative methods based on the relaxation time approximation (RTA) and iterative solutions of the Boltzmann transport equation (IBTE), as implemented in the EPW 46 code. Anharmonic phonons are obtained via SSCHA. The effect of local disorder is incorporated by substituting the electronic eigenvalues computed for the monomorphous structure with those obtained from the polymorphous structure, while retaining the wavefunctions of the monomorphous configuration. Despite this approximation, local disorder reduces the electron mobility by 43% at 300K, significantly improving agreement with experiment. An even greater reduction is achieved when local disorder is combined with electron-phonon coupling effects on the electronic structure, lowering the initial mobility overestimation from 838% to 290% 43, a remarkable improvement. Zhou and Bernardi 47 using the Perturbo code 48 achieved excellent agreement with experimental electron mobilities of SrTiO3 by combining the retarded cumulant approach with the Kubo formula, capturing both coherent quasiparticles and incoherent polaronic effects through the spectral function. While their method treats dynamical electron-phonon interactions accurately, it neglects local structural disorder. Combining these two effects remains an open and promising direction for future research, particularly in understanding their joint impact on charge transport 49. Poncé et al 44 investigated the origin of low carrier mobilities in halide perovskites by performing EPW calculations within the RTA for the orthorhombic phase of MAPbI3 and cubic phase of CsPbBr3. Their results, presented in Figure 4b, illustrate the influence of local disorder on carrier mobilities in an indirect yet insightful manner. Employing the orthorhombic structure, which incorporates the effect of octahedral tilting similar to local disorder in the cubic phase, yields better agreement with experimental data 50, 51. This holds even for temperatures above 300 K, where the tetragonal to cubic phase transition takes place. Instead, using monomorphous cubic CsPbI3 leads to a significant increase of the calculated carrier mobilities, leading to the overestimation of the experimental data by a factror of 7.
Figure 4c shows the calculated charge density of a large hole polaron in cubic CsPbBr3 using a $\times2$$\times$8 supercell. Polaron is a manifestation of electron-phonon coupling, where the carrier distorts the surrounding atomic structure and becomes dressed by phonons, forming a quasiparticle. The calculations were performed by removing one electron and performing a geometry relaxation using the hybrid PBE0 functional to explain large polarons observed experimentally using time-resolved optical Kerr effect spectroscopy 45. However, the computational approach employs a monomorphous cubic structure as the reference when introducing a charge and analyzing structural distortions. When a charged system is compared only to an idealized reference, the resulting lattice distortions may be overestimated or misattributed entirely to polaron formation. In reality, much of the distortion may already exist in the neutral, uncharged structure if modeled with a more physically accurate, symmetry-broken polymorphous configuration. Failing to use a locally disordered structure can lead to misleading conclusions about the nature, size, and stabilization energy of polarons in soft, anharmonic materials like CsPbBr3.
5 Future directions
The findings discussed in this Perspective establish the importance of incorporating local disorder through polymorphous structural models to accurately predict and understand the electronic, vibrational, and electron-phonon properties of soft, anharmonic high temperature phases of materials. Despite the progress, the theoretical treatment of electron-phonon coupling in the presence of positional polymorphism remains at an early stage, offering numerous opportunities for future research.
One promising avenue is the integration of polymorphous supercells with state-of-the-art methodologies for electron-phonon coupling, such as the use of perturbative approaches implemented in state-of-the-art codes like EPW 46 and Perturbo 48. These methods typically assume high symmetry reference structures, limiting their predictive power in materials where symmetry breaking is intrinsic. Developing workflows with locally disordered configurations could significantly improve accuracy in modeling carrier lifetimes, charge transport, and phonon-assisted optical absorption. Regarding charge transport, another promising direction is to advance non-perturbative supercell approaches for anharmonic electron-phonon coupling, allowing the evaluation of carrier mobilities that include temperature-dependent band gap renormalization on top of local disorder effects 49. An advantage of supercell approaches, despite lacking the elegance of fine Brillouin zone sampling, is their ability to seamlessly treat anharmonicity and electron-phonon coupling on the same footing, without relying on PES derivatives along harmonic phonons 47.
Another exciting direction lies in exploring non-equilibrium carrier dynamics in disordered lattices. First-principles simulations of carrier-phonon and phonon-phonon interactions in the ultrafast regime 52, can be extended to polymorphous structures to probe transient phenomena such as hot carrier relaxation, carrier localization, and coherent phonon generation, as well as to uncover scattering channels that are normally forbidden by symmetry in the monomorphous case. In parallel, excitonic effects in materials with strong local disorder remain largely unexplored; investigating how positional polymorphism influences exciton binding energies, lifetimes, and dissociation dynamics could offer new insights into light-matter interactions and optical response. Similarly, Auger recombination and radiative carrier lifetimes, which are central to optoelectronic device performance 53, can be re-evaluated.
The application of these frameworks to a broader class of materials, including layered/2D perovskites, double perovskites, antiperovskites, and quantum dots is another frontier. These systems often feature pronounced anisotropy and molecular complexity, making them ideal candidates for a polymorphous description. Particularly, the interplay of molecular cation dynamics with inorganic octahedral tilting and quantum confinement, deserves detailed investigation, potentially requiring the development of hybrid methodologies capable of treating the different sources of local disorder consistently 24.
It will also be interesting to compile databases of known and predicted anharmonic materials and conducting high-throughput first-principles studies to systematically explore their polymorphous structural networks across a broad chemical space 21. This effort would lay the groundwork for identifying trends and guiding the discovery of materials where local disorder plays a functional role. Moreover, given the large configurational space associated with polymorphous networks, machine learning force fields offer a powerful tool to accelerate structure exploration in large supercells and thus capture long range disorder as well as perform classification and even optoelectronic property prediction. This effort could also provide insight into how polar symmetry breaking is entangled with local disorder in complex material systems, including but not limited to Ge-based perovskites.
It is worth noting that the polymorphous framework has certain limitations. These include the lack of an explicit treatment of relaxation processes detectable by low-energy spectroscopies and the restricted supercell sizes, which may underestimate correlation lengths and produce simulated Bragg-like features instead of experimental diffuse scattering. Furthermore, in its current implementation, local disorder is explored at fixed lattice parameters, neglecting strain-structure coupling and potentially biasing the accessible minima. Future extensions should incorporate variable-cell relaxations and larger supercells. Another limitation of the polymorphous framework is its quasi-static, Born-Oppenheimer treatment, which neglects non-adiabatic forces which can govern photoinduced structural dynamics 54.
The full integration of the SCP theory with polymorphous structures offers a rigorous path toward modeling strongly anharmonic vibrations and demonstrate further the breakdown of the phonon quasiparticle picture. Such a framework would enable temperature-dependent phonon dispersions, lifetimes, thermal conductivity, and electron-phonon coupling coefficients to be computed on disordered environments, resolving current ambiguities in interpreting temperature-induced property changes. Modeling thermal conductivity in locally disordered materials is critical, since the mean free paths can be reduced down to interatomic distances 10, and scattering becomes dominated by diffuson-like modes. This opens a path toward predicting ultralow thermal conductivity in materials exhibiting entropy-driven structural complexity and glass-like behavior, despite their crystalline nature 38, 36, 19. Thermal conductivity relies on higher-order interatomic force constants, particularly third- and fourth-order terms governing phonon-phonon scattering, making machine learning force fields an essential tool for accelerating their computation in complex, disordered systems. Furthermore, extending local disorder to consider temperature-driven phase transitions and disorder-enabled functionalities such as ferroelectricity, spin polymorphism in paramagnetic phases, ionic transport, or polaron formation represents a rich area for computational development.
From the experimental perspective, advancing characterization methods that are sensitive to local disorder is equally essential. While PDF analysis remains a powerful tool to probe local structure, complementary techniques such as high-temperature Raman spectroscopy, nuclear magnetic resonance measurements, inelastic neutron scattering for phonon DOS, and diffuse scattering can offer critical insights into vibrational dynamics and structural fluctuations in disordered systems 6, 12, 11, 37. These methods provide additional observables that can be directly compared with first-principles predictions, allowing for more rigorous validation of polymorphous models, as well as elucidate long-range correlated disorder. Establishing such cross-validation pipelines between theory and experiment is key to uncovering the main mechanisms underpinning the efficiency of functional materials.
6 Acknowledgments
We acknowledge computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.